The summarizing phase of the accounting cycle ends with the preparation of the financial statements.

Select one:

True

False

Question 2

At the end of the month, Maria Clara Parlor's total Liabilities and Owner's Equity is P 250,000. One fifth of which is the liability. The parlor's total assets is

Select one:

a. Answer not given

b. P 250, 000

c. P 50, 000

d. P 200, 000

Question 3

A non-exchange business transaction does not involve changes in quantifiable monetary values.

Select one:

True

False

Question 4

The transaction: Mr. A invested P20, 000 cash and equipment worth P30, 000 in a laundry shop business, will result to an increase in assets and increase in capital for the business.

Select one:

True

False

Question 5

Assets are anything of economic value and owned by the business entity.

Select one:

True

False

Question 6

Mr. X's Capital account has a balance of P 100,000 on March 31, 2017. Withdrawals for personal expenses during April amounted to P 15,000. At the end of April, the Capital account has a balance of P135, 000. Mr. X's net income for April is

Select one:

a. P 35,000

b. P 50,000

c. P 20,000

d. P 85,000

Question 7

Maria Clara has the following transactions during the month of March: paid salaries of house helper, P 3,500; paid utility expenses of the parlor, P 2,500; at the end of month, hired a parlor helper with a monthly salary of P5, 000. The total expenditures that should be reported for the month of March is

Select one:

a. P 7,500

b. P 11,000

c. P 6,000

d. P 2,500

Question 8

The credit side of T-account is used to record increases in liability and capital.

Select one:

True

False

Question 9

The account to be credited is indented following the account debited and the corresponding amount is placed on the extreme right money column of the journal.

Select one:

True

False

Question 10

The right or claim of the owner over the assets of the firm is called liabilities.

Select one:

True

False
Answer #1

Answer:

1. The summarizing phase of the accounting cycle ends with the preparation of the financial statements.

True

2. At the end of the month, Maria Clara Parlor's Total Liabilities and Owner's Equity is P 250,000. One-fifth of which is the liability. The parlor's total assets are

b. P 250, 000. As Total Assets = Total Liabilities + Owners Equity.

3. A non-exchange business transaction does not involve changes in quantifiable monetary values.

False, A non-exchange transaction is one in which one party receives something of value without directly giving value in exchange. Non-exchange transactions are events that do not involve physical exchanges but where changes in monetary values are determinable, e.g. wear and tear of equipment, fire loss, typhoon loss, etc.

4. The transaction: Mr. A invested P20, 000 cash, and equipment worth P30, 000 in a laundry shop business, will result to an increase in assets and an increase in capital for the business.

True, As Increase in Asset Increase in Capital from the equation, Asset = Liability + Capital

5. Assets are anything of economic value and owned by the business entity.

True

6.Mr. X's Capital account has a balance of P 100,000 on March 31, 2017. Withdrawals for personal expenses during April amounted to P 15,000. At the end of April, the Capital account has a balance of P135, 000. Mr. X's net income for April is

Select one:

b. P 50,000,

Capital = 100000 -15000 + X = 135000

X = 135000+15000-10000

X = 150000-100000 = 50000

7. Maria Clara has the following transactions during the month of March: paid salaries of house helper, P 3,500; paid utility expenses of the parlor, P 2,500; at the end of month, hired a parlor helper with a monthly salary of P5, 000. The total expenditures that should be reported for the month of March is

c. P 6,000, Paid for March - 3500+2500 = 6000 ( March end salary for helper not to include as an expense incurred is not for the month of March - As per matching concept)

Question 8

The credit side of T-account is used to record increases in liability and capital.

True

Question 9

The account to be credited is indented following the account debited and the corresponding amount is placed on the extreme right money column of the journal.

True

Question 10

The right or claim of the owner over the assets of the firm is called liabilities.

True